### v1

@@ Abstract
In this paper, we study the Aubin property of the Karush-Kuhn-Tucker solution mapping for the nonlinear semidefinite programming (NLSDP) problem at a locally optimal solution. In the literature, it is known that the Aubin property implies the constraint nondegeneracy by Fusek [SIAM J. Optim. 23 (2013), pp. 1041-1061] and the second-order sufficient condition by Ding et al. [SIAM J. Optim. 27 (2017), pp. 67-90]. Based on the Mordukhovich criterion, here we further prove that the strong second-order sufficient condition is also necessary for the Aubin property to hold. Consequently, several equivalent conditions including the strong regularity are established for NLSDP’s Aubin property. Together with the recent progress made by Chen et al. on the equivalence between the Aubin property and the strong regularity for nonlinear second-order cone programming [arXiv:2406.13798v1 (2024)], this paper constitutes a significant step forward in characterizing the Aubin property for general non-polyhedral $C^{2}$-cone reducible constrained optimization problems.
@@ 1 Introduction
${\mathcal{L}}(x,y):=f(x)+\langle y,G(x)\rangle,\quad(x;y)\in{\mathcal{X}}\times{\mathcal{Y}},$ (1.2)
and the first-order optimality condition of (1.1) is characterized by the Karush-Kuhn-Tucker (KKT) system
@@
The canonically perturbed version [6, Section 5.1.3] of (1.1) is given by
$\min_{x\in{\mathcal{X}}}\,f(x)-\langle a,x\rangle\quad\mbox{s.t.}\quad G(x)-b\in{\mathcal{K}},$ (1.4)
@@
$a=\nabla_{x}{\mathcal{L}}(x,y)\quad\mbox{and}\quad y\in{\mathcal{N}}_{{\mathcal{K}}}(G(x)-b).$ (1.5)
@@
${{\mathbb{S}}}_{{\rm KKT}}(a,b):=\{(x;y)\mid a=\nabla_{x}{\mathcal{L}}(x,y),\ y\in{\mathcal{N}}_{{\mathcal{K}}}(G(x)-b)\}.$ (1.6)
@@
As a core research topic in perturbation analysis of optimization problems, how the solution set ${{\mathbb{S}}}_{{\rm KKT}}(a,b)$ of (1.5) varies along with $(a,b)$ around the origin has been studied for a long time [6, 13, 14, 19, 22, 32, 40]. In a landmark paper of Robinson [34], the definition of strong regularity was introduced to extend the inverse and implicit function theorems to generalized equations. Note that the KKT system (1.3) can be equivalently reformulated to the generalized equation
$0\in\Phi(x,y):=\begin{pmatrix}\nabla_{x}{\mathcal{L}}(x,y)\\ -G(x)\end{pmatrix}+\begin{pmatrix}0\\ {\mathcal{N}}_{{\mathcal{K}}}^{-1}(y)\end{pmatrix}.$ (1.7)
@@
Then, we say that $(\bar{x};\bar{y})$ is a strongly regular solution of generalized equation (1.7) (or the KKT system (1.3)) if the inverse of the set-valued mapping
$\overline{\Phi}(x,y):=\begin{pmatrix}\nabla_{xx}^{2}{\mathcal{L}}(\bar{x},\bar{y})(x-\bar{x})+\nabla G(\bar{x})(y-\bar{y})\\ -G(\bar{x})-G^{\prime}(\bar{x})(x-\bar{x})\end{pmatrix}+\begin{pmatrix}{\mathcal{N}}_{{\mathcal{X}}}(x)\\ {\mathcal{N}}_{{\mathcal{K}}}^{-1}(y)\end{pmatrix}$
@@
has a Lipschitz continuous single-valued localization around $(0;0)\in{\mathcal{X}}\times{\mathcal{Y}}$ for $(\bar{x};\bar{y})$, where $\nabla_{xx}^{2}{\mathcal{L}}$ is the partial Hessian of ${\mathcal{L}}$ with respect to $x$. According to [11], such a strong regularity condition is equivalent to the condition that the solution mapping ${{\mathbb{S}}}_{\rm KKT}$ has a Lipschitz continuous single-valued localization around $((0;0);(\bar{x};\bar{y}))$, or equivalently, the mapping $\Phi$ in (1.7) is strongly metrically regular at $(\bar{x};\bar{y})$ for $(0;0)$. Another significant yet less restrictive concept in studying the variation of ${{\mathbb{S}}}_{\rm KKT}$ with respect to the perturbation is the Aubin property (c.f. [40, Section 9F] or [13, Section 3E] for a systematic introduction), which was originally called the “pseudo-Lipschitzian” property by Aubin [1]. The Aubin property of ${{\mathbb{S}}}_{\rm KKT}$ holds at $(0;0)\in{\mathcal{X}}\times{\mathcal{Y}}$ for $(\bar{x};\bar{y})\in{\mathbb{S}}_{{\rm KKT}}(0,0)$ if there exist a constant $\kappa>0$ and open neighborhoods ${\mathcal{U}}$ of $(0;0)$ and ${\mathcal{V}}$ of $(\bar{x};\bar{y})$ such that
@@
${{\mathbb{S}}}_{\rm KKT}(a^{\prime},b^{\prime})\cap{\mathcal{V}}\subseteq{{\mathbb{S}}}_{\rm KKT}(a,b)+\kappa\|(a^{\prime},b^{\prime})-(a,b)\|\mathds{B}_{{\mathcal{X}}\times{\mathcal{Y}}}\ \forall\,(a;b),(a^{\prime};b^{\prime})\in{\mathcal{U}},$
@@
where $\mathds{B}_{{\mathcal{X}}\times{\mathcal{Y}}}$ denotes the closed unit ball in ${{\mathcal{X}}\times{\mathcal{Y}}}$ centered at the origin. It is easy to see from definition that the Aubin property of ${{\mathbb{S}}}_{\rm KKT}$ holds at $(0;0)$ for $(\bar{x};\bar{y})$ if $(\bar{x};\bar{y})$ is a strongly regular solution to (1.3). Moreover, such an Aubin property is equivalent to the metric regularity or the linear openness of $\Phi$ at $(\bar{x},\bar{y})$ (e.g. [40, Theorem 9.43]).
@@
Since the solution mapping ${\mathbb{S}}_{{\rm KKT}}$ is implicitly defined, verifying both the strong regularity and the Aubin property from their definitions is generally unachievable. Consequently, equivalent characterizations for the two conditions have evolved into a central topic in optimization theory and variational analysis. For conventional nonlinear programming with ${\mathcal{K}}$ in (1.1) being a convex polyhedral cone, such equivalent characterizations have been well established for about three decades. Specifically, Robinson [34] defined the strong second-order sufficient condition (SSOSC) for the nonlinear programming problem. Moreover, at a stationary point $\bar{x}$ satisfying the linear independence constraint qualification (LICQ), he also showed in [34, Theorem 4.1] that the KKT system is strongly regular at $(\bar{x};\bar{y})\in{\mathbb{S}}_{\rm KKT}$ if the SSOSC holds at $(\bar{x};\bar{y})$. At the same time, Kojima [25] introduced the concept of strong stability to stationary points of the nonlinear programming problem, and showed in [25, Theorems 6.4 & 6.5] that for a locally optimal solution satisfying the LICQ, it is strongly stable if and only if the SSOSC holds. It was later revealed by Jongen et al. [20, Section 3] that the strong regularity and the strong stability are equivalent for stationary points of nonlinear programming where the LICQ holds. Furthermore, the strong regularity was characterized by Kummer [26] based on a generalized “implicit function theorem” on nonsmooth equations, and one may refer to [21] and [24] for a related approach, especially [24, Theorem 4.3] for a survey of equivalent characterizations. By combining the results of [34], [25], and [24], one has that at a locally optimal solution of the nonlinear programming problem, the strong regularity is equivalent to the condition that both the SSOSC and the LICQ hold (c.f. [5, Remark 4.11]). Such a result is also available in [5, Theorem 4.10] and [6, Proposition 5.38]. In addition to these equivalent characterizations of the strong regularity, a surprising result of Dontchev and Rockafellar [12, Theorems 1, 4, & 5] for the nonlinear programming problem is that the Aubin property of ${{\mathbb{S}}}_{\rm KKT}$ holds at $(0;0)$ for $(\bar{x};\bar{y})\in{\mathbb{S}}_{\rm KKT}$ with $\bar{x}$ being a locally optimal solution can imply that the strong regularity of the KKT system (1.3) at $(\bar{x};\bar{y})$. Consequently, a comprehensive class of equivalent characterizations of the Aubin property for the nonlinear programming problem was achieved.
@@
When the set ${\mathcal{K}}$ in (1.1) is no longer a polyhedral set, characterizing the two concepts becomes much more involved. A more realistic setting is that ${\mathcal{K}}$ is $C^{2}$-cone reducible (c.f. [6, Definition 3.135]), which is practical enough for encompassing many important classes of optimization problems, including the nonlinear programming, the nonlinear second-order cone programming (NLSOCP), and the nonlinear semidefinite programming (NLSDP) [41]. In this setting, the perturbation analysis of problem (1.1) has been extensively studied [2, 3, 4, 10, 42, 43], and characterizing the strong regularity via second-order optimality conditions has been recognized as a prominent topic. Although deriving a result akin to [6, Proposition 5.38] for problem (1.1) in its general form has not been achieved, achievements have been made for the most representative classes of problems with significant importance in the form of (1.1).
@@
According to [6, Theorem 5.24], the strong regularity of the KKT system (1.3) at a solution $(\bar{x};\bar{y})$ implies that the constraint nondegeneracy condition holds at $\bar{x}$ (or $\bar{x}$ is nondegenerate, c.f. (2.13)). With the help of this result, Bonans and Ramírez [4] established a counterpart of [6, Proposition 5.38] for the NLSOCP problem. For the NLSDP problem, Sun [42] defined the SSOSC by introducing an approximation set, and finally obtained a collection of equivalent characterizations of the strong regularity condition, including the SSOSC accompanied by the constraint nondegeneracy. These results properly extend the characterizations of the strong regularity from the conventional nonlinear programming to problem (1.1) with ${\mathcal{K}}$ being a non-polyhedral set. Nevertheless, for both the NLSOCP and the NLSDP problems, obtaining such an extension for characterizing the Aubin property has been an open question for a long time.
@@
The first step in addressing this issue was achieved by Outrata and Ramírez [31] (and the erratum by Opazo, Outrata, and Ramírez [30]). They proved that, for a nondegenerate locally optimal solution $\bar{x}$ of the NLSOCP problem (i.e., (1.1) with ${\mathcal{K}}$ being the Cartesian product of second-order cones), the Aubin property of a solution mapping (akin to (3.1)) at $0\in{\mathcal{X}}$ for $\bar{x}$ can imply the SSOSC, but under an assumption regarding the strict complementarity. Recently, Chen et al. [7] finalized this conclusion by removing these assumptions, and showed that, with $\bar{y}\in{\mathcal{M}}(\bar{x})$, the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ at $(0;0)$ for $(\bar{x};\bar{y})$ is equivalent to the strong regularity of $(\bar{x};\bar{y})$ to the KKT system, constituting a counterpart of the seminal result of Dontchev and Rockafellar [12] for the NLSOCP. As a result, the extension of the characterizations for the Aubin property is realized from the nonlinear programming problem to the NLSOCP problem.
@@
In this paper, we study the equivalent characterizations of the Aubin property for the NLSDP problem
@@
where $h:{\mathcal{X}}\to{\mathds{R}}^{m}$ and $g_{j}:{\mathcal{X}}\to{\mathcal{S}}^{p_{j}}_{+}$, $j=1,\ldots,J$ are twice continuously differentiable functions, ${\mathcal{S}}^{p_{j}}$ is the linear space of ${p_{j}}\times{p_{j}}$ real symmetric matrices endowed with the inner product $\langle A,B\rangle:={\sf trace}(AB)$ for $A,B\in{\mathcal{S}}^{p_{j}}$, where ${\sf trace}(\cdot)$ denotes the sum of the diagonal elements, and ${\mathcal{S}}^{p_{j}}_{+}$ is the closed convex cone of all positive semidefinite matrices in ${\mathcal{S}}^{p_{j}}$. When $J=1$ in (1.8), the strong regularity of the corresponding KKT system is equivalent to the SSOSC together with the constraint nondegeneracy [42, Theorem 4.1]. For the case that $J>1$, it is easy to see from the analysis in [42] that such an equivalence still holds. Moreover, in the case that $J=1$, Fusek [15] showed that the Aubin property at $(0;0)$ for a solution $(\bar{x};\bar{y})$ can imply the constraint nondegeneracy at $\bar{x}$. Such a result was later extended by Klatte and Kummer [23] to the general case of problem (1.1).
The main contribution of this paper is that, at a locally optimal solution $\bar{x}$ of (1.8) with $\bar{y}\in{\mathcal{M}}(\bar{x})$, we prove that the Aubin property of ${\mathbb{S}}_{\rm KKT}$ at $(0;0)$ for $(\bar{x};\bar{y})$ implies the SSOSC. We achieve this by designing an auxiliary optimization problem and fully exploiting its properties, and using the Mordukhovich criterion for characterizing the Aubin property [27]. We should emphasize that the key tools we used in this paper are essentially different from those in [7] for the nonlinear second-order cone programming. Specifically, the main progress in [7] is based on a lemma of alternative choices on cones, and the fact that the spectral factorization associated with second-order cones admits only two “eigenvalues” is indispensable. This approach can be used here when $\max_{j}\{p_{j}\}\leq 3$, but does not apply to the general cases. Consequently, it is not hard to see that the tools in [7] are far from sufficient for deriving a counterpart of its main result in the setting of the NLSDP problem (1.8). Based on the equivalence between the Aubin property and the strong regularity established in this paper, we finally obtain a comprehensive set of equivalent characterizations of the Aubin property for the NLSDP problem (1.8).
The remaining parts of this paper are organized as follows. Section 2 introduces the notation, the definitions, and the preliminary results used throughout this paper. In Section 3, we study the implications of the Aubin property of the solution mapping ${\mathbb{S}}_{{\rm KKT}}(a,b)$, especially the SSOSC. The equivalent characterizations of the Aubin property are given in Section 4. We conclude this paper in Section 5.

### v2

@@ Abstract
In this paper, we study the Aubin property of the Karush-Kuhn-Tucker solution mapping for the nonlinear semidefinite programming (NLSDP) problem at a locally optimal solution. In the literature, it is known that the Aubin property implies the constraint nondegeneracy by Fusek [SIAM J. Optim. 23 (2013), pp. 1041-1061] and the second-order sufficient condition by Ding et al. [SIAM J. Optim. 27 (2017), pp. 67-90]. Based on the Mordukhovich criterion, here we further prove that the strong second-order sufficient condition is also necessary for the Aubin property to hold. Consequently, several equivalent conditions including the strong regularity are established for NLSDP’s Aubin property. Together with the recent progress made by Chen et al. on the equivalence between the Aubin property and the strong regularity for nonlinear second-order cone programming [SIAM J. Optim., in press; arXiv:2406.13798v3 (2024)], this paper constitutes a significant step forward in characterizing the Aubin property for general non-polyhedral $C^{2}$-cone reducible constrained optimization problems.
@@ 1 Introduction
${\mathcal{L}}(x,y):=f(x)+\langle y,G(x)\rangle,\quad(x;y)\in{\mathcal{X}}\times{\mathcal{Y}}.$ (1.2)
Here and throughout this paper, the notation (;)($\cdot$;$\cdot$)( ⋅ ; ⋅ ) signifies the arrangement of two vectors or linear operators symbolically column-wise. Then the first-order optimality condition of (1.1) is characterized by the Karush-Kuhn-Tucker (KKT) system
@@
The canonically perturbed version [7, Section 5.1.3] of (1.1) is given by
$\min_{x\in{\mathcal{X}}}\,f(x)-\langle a,x\rangle\quad\mbox{s.t.}\quad G(x)+b\in{\mathcal{K}},$ (1.4)
@@
$a=\nabla_{x}{\mathcal{L}}(x,y)\quad\mbox{and}\quad y\in{\mathcal{N}}_{{\mathcal{K}}}(G(x)+b).$ (1.5)
@@
${{\mathbb{S}}}_{{\rm KKT}}(a,b):=\{(x;y)\mid a=\nabla_{x}{\mathcal{L}}(x,y),\ y\in{\mathcal{N}}_{{\mathcal{K}}}(G(x)+b)\}.$ (1.6)
@@
As a core research topic in perturbation analysis of optimization problems, how the solution set ${{\mathbb{S}}}_{{\rm KKT}}(a,b)$ of (1.5) varies along with $(a;b)$ around the origin has been studied for a long time [7, 14, 15, 21, 24, 34, 43]. In a landmark paper of Robinson [37], the definition of strong regularity was introduced to extend the inverse and implicit function theorems to generalized equations. Note that the KKT system (1.3) can be equivalently reformulated to the generalized equation
$0\in\Phi(x,y):=\begin{pmatrix}\nabla_{x}{\mathcal{L}}(x,y)\\ -G(x)\end{pmatrix}+\begin{pmatrix}\{0\}\\ {\mathcal{N}}_{{\mathcal{K}}}^{-1}(y)\end{pmatrix}.$ (1.7)
@@
Then, we say that $(\bar{x};\bar{y})\in{\mathcal{X}}\times{\mathcal{Y}}$ is a strongly regular solution of (1.7) (or the KKT system (1.3)) if the inverse of the set-valued mapping
$\overline{\Phi}(x,y):=\begin{pmatrix}\nabla_{xx}^{2}{\mathcal{L}}(\bar{x},\bar{y})(x-\bar{x})+\nabla G(\bar{x})(y-\bar{y})\\ -G(\bar{x})-G^{\prime}(\bar{x})(x-\bar{x})\end{pmatrix}+\begin{pmatrix}\{0\}\\ {\mathcal{N}}_{{\mathcal{K}}}^{-1}(y)\end{pmatrix}$
@@
has a Lipschitz continuous single-valued localization around $(0;0)\in{\mathcal{X}}\times{\mathcal{Y}}$ for $(\bar{x};\bar{y})$, where $\nabla_{xx}^{2}{\mathcal{L}}$ is the partial Hessian of ${\mathcal{L}}$ with respect to $x$. According to [12], such a strong regularity condition is equivalent to the condition that the solution mapping ${{\mathbb{S}}}_{\rm KKT}$ has a Lipschitz continuous single-valued localization around $((0;0);(\bar{x};\bar{y}))$, or equivalently, the mapping $\Phi$ in (1.7) is strongly metrically regular at $(\bar{x};\bar{y})$ for $(0;0)$. Another significant yet less restrictive concept in studying the variation of ${{\mathbb{S}}}_{\rm KKT}(a,b)$ with respect to the perturbation is the Aubin property (cf. [43, Section 9F] or [14, Section 3E] for a systematic introduction), which was originally called the “pseudo-Lipschitzian” property by Aubin [1]. The Aubin property of ${{\mathbb{S}}}_{\rm KKT}$ holds at $(0;0)\in{\mathcal{X}}\times{\mathcal{Y}}$ for $(\bar{x};\bar{y})\in{\mathbb{S}}_{{\rm KKT}}(0,0)$ if there exist a constant $\kappa>0$ and open neighborhoods ${\mathcal{U}}$ of $(0;0)$ and ${\mathcal{V}}$ of $(\bar{x};\bar{y})$ such that
@@
${{\mathbb{S}}}_{\rm KKT}(a^{\prime},b^{\prime})\cap{\mathcal{V}}\subseteq{{\mathbb{S}}}_{\rm KKT}(a,b)+\kappa\|(a^{\prime};b^{\prime})-(a;b)\|\mathds{B}_{{\mathcal{X}}\times{\mathcal{Y}}}\ \,\forall\,(a;b),(a^{\prime};b^{\prime})\in{\mathcal{U}},$
@@
where $\mathds{B}_{{\mathcal{X}}\times{\mathcal{Y}}}$ denotes the closed unit ball in ${{\mathcal{X}}\times{\mathcal{Y}}}$ centered at the origin. It is easy to see from the definition that the Aubin property of ${{\mathbb{S}}}_{\rm KKT}$ holds at $(0;0)$ for $(\bar{x};\bar{y})$ if $(\bar{x};\bar{y})$ is a strongly regular solution to (1.3). Moreover, such an Aubin property is equivalent to the metric regularity or the linear openness of $\Phi$ at $(\bar{x};\bar{y})$ (e.g. [43, Theorem 9.43]).
@@
Since the solution mapping ${\mathbb{S}}_{{\rm KKT}}$ is implicitly defined, verifying both the strong regularity and the Aubin property from their definitions is generally unachievable. Consequently, characterizations for the two conditions have evolved into a central topic in optimization theory and variational analysis. For conventional nonlinear programming with ${\mathcal{K}}$ in (1.1) being a convex polyhedral cone, such characterizations have been well established for about three decades. Specifically, Robinson [37] defined the strong second-order sufficient condition (SSOSC) for the nonlinear programming problem. Moreover, at a stationary point $\bar{x}$ satisfying the linear independence constraint qualification (LICQ), he also showed in [37, Theorem 4.1] that the KKT system is strongly regular at $(\bar{x};\bar{y})\in{\mathbb{S}}_{\rm KKT}$ if the SSOSC holds at $(\bar{x};\bar{y})$. At the same time, Kojima [27] introduced the concept of strong stability to stationary points of the nonlinear programming problem, and showed in [27, Theorems 6.4 & 6.5] that for a locally optimal solution satisfying the LICQ, it is strongly stable if and only if the SSOSC holds. It was later revealed by Jongen et al. [23, Section 3] that the strong regularity and the strong stability are equivalent for stationary points of nonlinear programming where the LICQ holds. Furthermore, the strong regularity was characterized by Kummer [28] based on a generalized “implicit function theorem” on nonsmooth equations, and one may refer to [22] and [26] for a related approach, especially [26, Theorem 4.3] for a survey of equivalent characterizations. By combining the results of [37], [27], and [26], one has that at a locally optimal solution of the nonlinear programming problem, the strong regularity is equivalent to the condition that both the SSOSC and the LICQ hold (cf. [6, Remark 4.11]). Such a result is also available in [6, Theorem 4.10] and [7, Proposition 5.38]. In addition to these equivalent characterizations of the strong regularity, a surprising result of Dontchev and Rockafellar [13, Theorems 1, 4, & 5] for the nonlinear programming problem is that the Aubin property of ${{\mathbb{S}}}_{\rm KKT}$ at $(0;0)$ for $(\bar{x};\bar{y})\in{\mathbb{S}}_{\rm KKT}(0,0)$, with $\bar{x}$ being a locally optimal solution, implies the strong regularity of the KKT system (1.3) at $(\bar{x};\bar{y})$. Consequently, a comprehensive class of equivalent characterizations of the Aubin property for the nonlinear programming problem was achieved.
@@
When the set ${\mathcal{K}}$ in (1.1) is no longer a polyhedral set, characterizing these two concepts becomes much more involved. A more realistic setting is that ${\mathcal{K}}$ is $C^{2}$-cone reducible (cf. [7, Definition 3.135]), which is practical enough for encompassing many important classes of optimization problems, including the nonlinear programming, the nonlinear second-order cone programming (NLSOCP), and the nonlinear semidefinite programming (NLSDP) [44]. In this setting, the perturbation analysis of problem (1.1) has been extensively studied [3, 4, 5, 11, 45, 46], and characterizing the strong regularity via second-order optimality conditions has been recognized as a prominent topic. Although deriving an equivalent characterization of the strong regularity akin to [7, Proposition 5.38] for problem (1.1) in its general form has not been achieved, achievements have been made for the most representative classes of problems with significant importance in the form of (1.1).
@@
According to [7, Theorem 5.24], the strong regularity of the KKT system (1.3) at a solution $(\bar{x};\bar{y})$ implies that the constraint nondegeneracy condition holds at $\bar{x}$ (or $\bar{x}$ is nondegenerate, cf. (2.13)). With the help of this result, Bonnans and Ramírez [5] established a counterpart of [7, Proposition 5.38] for the NLSOCP problem. For the NLSDP problem, Sun [45] defined the SSOSC by introducing an approximation set, and finally obtained a collection of equivalent characterizations of the strong regularity condition, including the SSOSC accompanied by the constraint nondegeneracy. These results properly extend the characterizations of the strong regularity from the conventional nonlinear programming to problem (1.1) with ${\mathcal{K}}$ being a non-polyhedral set. Nevertheless, for both the NLSOCP and the NLSDP problems, obtaining such an extension for characterizing the Aubin property has been an open question for a long time.
@@
The first step to address this issue was achieved by Outrata and Ramírez [33] (and the erratum by Opazo, Outrata, and Ramírez [32]). They proved that, for a nondegenerate locally optimal solution $\bar{x}$ of the NLSOCP problem (i.e., (1.1) with ${\mathcal{K}}$ being the Cartesian product of second-order cones), the Aubin property of a solution mapping (akin to (3.1)) at $0\in{\mathcal{X}}$ for $\bar{x}$ implies the SSOSC, but under an assumption regarding the strict complementarity. Recently, Chen et al. [8] finalized this conclusion by removing these assumptions and showed that, with $\bar{y}\in{\mathcal{M}}(\bar{x})$, the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ at $(0;0)$ for $(\bar{x};\bar{y})$ is equivalent to the strong regularity of $(\bar{x};\bar{y})$ to the KKT system, thus constituting for the NLSOCP a counterpart of the seminal result of Dontchev and Rockafellar [13]. As a result, an extension of the characterizations for the Aubin property is realized from the nonlinear programming problem to the NLSOCP problem. Generally, if ${\mathcal{K}}$ is a $C^{2}$-cone reducible set, another recent work [19, Theorem 5.14] utilized an assumption on relative interiors of subdifferential mappings to prove the equivalence between the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ and the strong regularity of the KKT system, but such an assumption is exactly the strict complementarity condition when applied to the desired NLSDP problem.
@@
In this paper, we study the characterizations of the Aubin property for the KKT solution mapping of the NLSDP problem
@@
where $h:{\mathcal{X}}\to{\mathds{R}}^{m}$ and $g_{j}:{\mathcal{X}}\to{\mathcal{S}}^{p_{j}}$, $j=1,\ldots,J$ are twice continuously differentiable functions, ${\mathcal{S}}^{p_{j}}$ is the linear space of ${p_{j}}\times{p_{j}}$ real symmetric matrices endowed with the inner product $\langle A,B\rangle:={\sf trace}(AB)$ for $A,B\in{\mathcal{S}}^{p_{j}}$, where ${\sf trace}(\cdot)$ denotes the sum of the diagonal elements, and ${\mathcal{S}}^{p_{j}}_{+}$ is the closed convex cone of all positive semidefinite matrices in ${\mathcal{S}}^{p_{j}}$. When $J=1$ in (1.8), the strong regularity of the corresponding KKT system is equivalent to the SSOSC together with the constraint nondegeneracy [45, Theorem 4.1]. For the case that $J>1$, it is easy to see from the analysis in [45] that such an equivalence still holds. Moreover, in the case that $J=1$, Fusek [16] showed that the Aubin property at $(0;0)$ for a solution $(\bar{x};\bar{y})$ implies the constraint nondegeneracy at $\bar{x}$. Such a result was later extended by Klatte and Kummer [25] to the general case of problem (1.1).
The main contribution of this paper is that, at a locally optimal solution $\bar{x}$ of (1.8) with $\bar{y}\in{\mathcal{M}}(\bar{x})$, we prove that the Aubin property of ${\mathbb{S}}_{\rm KKT}$ at $(0;0)$ for $(\bar{x};\bar{y})$ implies the SSOSC. We achieve this by designing an auxiliary optimization problem and fully exploiting its properties, and using the Mordukhovich criterion for characterizing the Aubin property [29]. We should emphasize that the key tools we used in this paper are essentially different from those in [8] for the NLSOCP. Specifically, the main progress in [8] is based on a lemma of alternative choices on cones, and the fact that the spectral factorization associated with second-order cones admits only two “eigenvalues” is indispensable. This approach can be used here when $\max_{j}\{p_{j}\}\leq 3$, but does not apply to the general cases. Consequently, it is not hard to see that the tools in [8] are far from sufficient for deriving a counterpart of its main result in the setting of the NLSDP problem (1.8). Based on the equivalence between the Aubin property and the strong regularity established in this paper, we finally obtain a comprehensive collection of equivalent characterizations of the Aubin property for the NLSDP problem (1.8).
The remaining parts of this paper are organized as follows. Section 2 introduces the notation, the definitions, and the preliminary results used throughout this paper. In Section 3, we study the implications of the Aubin property of the solution mapping ${\mathbb{S}}_{{\rm KKT}}$ for NLSDP, especially the SSOSC. The characterizations of the Aubin property of ${\mathbb{S}}_{{\rm KKT}}$ for NLSDP are given in Section 4. We conclude this paper with some discussions in Section 5.